A 55.0 kg player vaults to a height of 6.5 m before dropping to a think foam mattress that was placed to cushion her fall. Find (a) the momentum of the player just before landing and (b) the force exerted by the mattress if the player was stopped in 0.65s.
1
Expert's answer
2020-12-03T06:52:40-0500
"v=\sqrt{2gh}\\p=mv=m\sqrt{2gh}\\p=55\sqrt{2(9.8)(6.5)}=620\frac{kg\cdot m}{s}"

"F=\frac{p}{t}=\frac{620}{0.65}=955\ N"